A new security system needs to be evaluated in the airport. The probability of a person being a security hazard is 4%. At the checkpoint, the security system denied a person without security problems 1.5% of the time. Also the security system passed a person with security problems 1% of the time.

What is the probability that a random person does not pass through the system and is without any security problems?
Answer #1

P( a random person does not pass through the system and is without any security problems )

=P(not a security hazard)*P(not pass through| not a security hazard)

=(1-0.04)*0.015=0.0144